OpenStudy (anonymous):

what is the of a parabola inscribed in rectangle 30cm long and 22cm wide?

OpenStudy (ybarrap):

x^2=4py At x=22/2 = 11 , y = 30 So p = x^2/4y = 11^2/4(30)=121/120 (btw, this is the focus) So you equation of a parabola is y = x^2 / 4 * p = x^2 /(4*121/120) = 30x^2/121 http://www.wolframalpha.com/input/?i=plot+30*x%5E2%2F121%2C+-11%3C%3Dx%3C%3D11

OpenStudy (anonymous):

may I know what formula did u use?

OpenStudy (ybarrap):

of course - http://en.wikipedia.org/wiki/Parabola Since this is centered at the origin and symmetric about the y-axis, we use x^2=4py. Does this make sense?